Using CAPM determine the required rate of return on the following:

            a.         Rf = 3.10%                   b.         Rf = 5.95%                   c.         Rf = 2.29%

                        Rm = 10.12%                           Rm = 12.37%                           Rm = 8.58%

                        b= 1.21                                   b= 0.78                                   b= 1.55
Answer #1

The required return on a stock is calculated using the Capital Asset Pricing Model (CAPM)

The formula is given below:

Ke=Rf+b[E(Rm)-Rf]

where:

Rf=risk-free rate of return which is the yield on default free debt like treasury notes

Rm=expected rate of return on the market.

b= Stock’s beta

1.Ke= 3.10 + 1.2*(10.12 – 3.10)

        = 3.10 + 1.2*7.02

        = 3.10 + 8.42

        = 11.52%.

Therefore, the required return is 11.52%.

2.Ke= 5.95 + 0.78(12.37 – 5.95)

        = 5.95 + 0.78*6.42

        = 5.95 + 5.01

        = 19.96.

Therefore, the required return is 19.96%.

3.Ke= 2.99 + 1.55*(8.58 – 2.99)

        = 2.99 + 1.55*5.59

        = 2.99 + 10.06

        = 13.05.

Therefore, the required return is 13.05%.

In case of any query, kindly comment on the solution.